OpenStudy (anonymous):

if a number n is added to the numerator 7/8 and twice as much is added to the denominator the result is 3/4. find the number.

OpenStudy (anonymous):

So, we know that \[\frac{ 7 + n }{ 8 + 2n} = \frac{ 3 }{ 4 }\] Solve for n to get n = 2.

OpenStudy (anonymous):

thanks